KyledaGreat:

Determine h(x^2) for the following function.

KyledaGreat:

\[h(x) = \frac{ -4x }{ x^2 - 9x - 20}\]

KyledaGreat:

\[h \left( x^{2}\right) = \]

Tranquility:

Everywhere you see x, replace it with x^2

Tranquility:

\(h(\color{red}{x}) = \dfrac{ -4\color{red}{x}}{\color{red}{ x}^2 - 9\color{red}{x} - 20}\) \(h(\color{red}{x^2}) = ~?\)

KyledaGreat:

h(x^2) = -4x^2/x^2 - 9x^2 - 20 is that what you mean

Tranquility:

Almost... but when you replace the x in x^2 with x^2 It would become (x^2)^2 which is equal to x^(2*2) = x^4

KyledaGreat:

x^4

Tranquility:

Yes. Can you write the full complete answer now that you know it's supposed to be x^4 \( (x^2)^2 = x^{2\cdot2} = x^4\)

KyledaGreat:

h(x^2) = x^4

Tranquility:

no...

Tranquility:

\(h(\color{red}{x}) = \dfrac{ -4\color{red}{x}}{\color{red}{ x}^2 - 9\color{red}{x} - 20}\) \(h(\color{blue}{x^2}) = \dfrac{ -4\color{blue}{x^2}}{\color{blue}{ (x^2)}^2 - 9\color{blue}{x^2} - 20}\)

KyledaGreat:

oh okay i see, i'm sorry

Tranquility:

No worries, as long as you understand now!